OpenStudy (anonymous):

The sum of three consecutive integers is 47 is less than the least of the integers. Find the integers.

hartnn (hartnn):

can u check/ verify your question ?

OpenStudy (anonymous):

The sum of three consecutive integers is 47 less than the least of the integers. Find the integers.

hartnn (hartnn):

yes, now it makes sense.

hartnn (hartnn):

let the least integer be x then next two integers will be x+1 and x+2 did u understand this ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

would it be 3x+3=x-47?

hartnn (hartnn):

so since the sum is 47 less than x, sum = x+x+1+x+2 =x-47 yes, thats correct.

OpenStudy (anonymous):

thanks. i got it

hartnn (hartnn):

welcome :)